# MM-Interleaved: Interleaved Image-Text Generative Modeling via Multi-modal Feature Synchronizer

### v1

@@ 4.2 Evaluation
(a) pre-training with 224 input resolution
@@
(b) fine-tuning with 448 input resolution
@@
Table 5: Ablation on the usage of MMFS module. “Genetation” is the text-to-image generation task. “ADE20k” is the segmentation-to-image translation task. Others are text generation tasks. “# Token” indicates the number of input visual tokens for LLM (32 by default).
@@
Table 6: Ablation on the design choice of MMFS module. Different attention modules can be used in MMFS. We also ablate whether to add additional transition layer before feeding image features into MMFS. Transformer attention with Resampler transition on single scale and single image is similar to the cross attention used in Flamingo. “Self-Attention” has the same parameters with “Resampler”.
4.3 Ablation Study
In the ablation study, we use CLIP-ViT-L/14 [70] as the vision foundation model, OpenLLaMA-3B v2 [26] as the multi-modal LLM and miniSD [^1] https://huggingface.co/justinpinkney/miniSD as the image decoder. The output resolution of each image is 256.
For simplicity, in this subsection, we study the effect of token efficiency, image resolution, cross-attention mechanisms, and the role of MMFS in image generation in a single-image setting, where each token only attends to the nearest preceding image. Please refer to the appendix for more detailed ablations studies in the multi-image setting.
The model is pre-trained for 10k steps on 80 NVIDIA 80G-A100 GPUs. The mini batch size of LAION-COCO [77], LAION-2B [76], and MMC4 [109] on a single GPU is 2, 2, and 4 respectively. For token efficiency, image resolution, and cross-attention mechanism, we evaluate the model’s zero-shot performance on three representative tasks and four datasets, i.e., image caption on the COCO Karpathy test set, text-to-image generation on the COCO Karpathy test set, visual question answering on OKVQA [59], and TextVQA validation set. We use CIDEr, FID-5k, and top-1 accuracy as the evaluation metrics. For cross-attention in image generation, we perform supervised fine-tuning on the ADE20K [106] for the label-to-image generation task.
Token Efficiency. As shown in Tab. 4(a), our method achieves better overall performance while only using 32 visual tokens for each input image. Such results demonstrate the effectiveness of our method when the context length is limited. To avoid the potential information bottleneck caused by the Resampler, we further increase the number of visual tokens to 256. Our method still outperforms the baseline in this setting.
@@ 4.3 Ablation Study
Scaling up Input Image Resolutions. The performance gap between our method and the baseline becomes larger when further increasing the input image resolution from 224 to 448, as shown in Tab. 4(b). Such results indicate our method could better exploit the additional information gained from high resolution.
Comparison between Different Cross Attention Mechanisms. The ablation of adopting different cross-attention mechanisms for LLM is shown in Tab. 6. The overall performance drops when directly replacing MMFS with the vanilla dense cross attention, possibly due to its slower convergence speed on 2D data, and using a resampler performs equally well. We highlight that the model with deformable attention performs significantly better than other attention mechanisms on TextVQA, indicating that deformable attention could effectively and efficiently capture fine-grained information like text needed for the task, such as visual question answering in this case.
@@
MMFS for Image Generation. We study whether adding MMFS is critical for image generation on the segmentation-to-image task. This task is hard as it requires precious pixel-level information to align the given segmentation condition and image output properly. From the results in Tab. 4(c), it is clear that the model without MMFS failed to generate images spatially aligned with the segmentation image condition, showing extremely low mIoU results. We show visualizations of the generated results in the supplementary material to provide a visual comparison. It is worth noting that the quality of the generated image from the model without MMFS is still high and properly aligned with the text description input; however, without MMFS, the model cannot preserve all spatial information, and the spatial alignment for the generated results is poor.
5 Conclusion
In conclusion, this paper presented MM-Interleaved, an end-to-end generative model specifically designed for interleaved image-text data. The unique contribution of MM-Interleaved lies in the novel multi-scale and multi-image feature synchronizer module, which significantly enhances the model’s capability to access and integrate fine-grained image features. This design addresses the longstanding limitation in existing models where a fixed number of visual tokens inadequately capture intricate image details, especially in scenarios involving multiple images. Pre-trained on diverse image-text datasets and further refined through supervised fine-tuning, MM-Interleaved shows a respectful proficiency in processing complex multi-modal instructions. The empirical results modestly highlight its ability to discern intricate visual details and generate images that align accurately with both textual and visual inputs.
Appendix A Additional Ablation Studies
Pre-training with Different Loss Terms. As showed in Tab. 7, MM-Interleaved achieves better performance when jointly trained with both $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ on comprehension and generation tasks, demonstrating the mutual benefits between the two loss terms. Moreover, setting $\lambda=10$ achieves a better balance between $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ empirically.
[…]
$\mathcal{L}_{NTP}+100~{}\mathcal{L}_{NIP}$ 106.2 31.1 29.8 24.5
$\mathcal{L}_{NTP}+10~{}\mathcal{L}_{NIP}$ 110.6 30.0 29.8 27.7
$\mathcal{L}_{NTP}+\mathcal{L}_{NIP}$ 110.0 31.4 29.3 26.0
$\mathcal{L}_{NTP}$ only 105.7 29.9 27.6
$\mathcal{L}_{NIP}$ only 34.2
Table 7: Pre-training with different loss terms.
[…]
110.6 30.0 29.8 27.7
107.0 32.2 28.7 22.5
102.7 32.0 27.3 22.0
MMFS with Multi-Image and Multi-Scale. As shown in Tab. 9, adding multi-image and multi-scale for MMFS improves the model’s performance on VQA tasks. To better take advantage of the multi-image mechanism, we follow [3, 83] to further evaluate our model with few-shot prompts. Two training samples including both image and text are selected as the input prompts. The results in Tab. 10 demonstrates that MMFS with multi-image and multi-scale benefits the most from the few-shot setting, indicating its potential capability for interleaved image-text modeling.
@@ Appendix A Additional Ablation Studies
Table 9: Zero-shot evaluation results of MMFS.
[…]
36.5 30.7
38.5 31.1
38.7 33.1
finetuning with 448 input resolution
38.3 36.6
38.7 37.7
39.1 37.9
Table 10: 2-shot evaluation results of MMFS on VQA tasks.
Appendix B Implementation Details
B.1 Pre-training
@@
Data Concatenation Strategy. For image-text-pair datasets (e.g., LAION-2B, LAION-COCO [77]), we randomly sample multiple image-text pairs from the same dataset and concatenate them to the maximum context length (i.e., 2048) during pre-training. For interleaved image and text datasets (i.e., MMC4 [109]), we also split and concatenate the documents to form the training samples. Such concatenation strategy can utilize the full context window of Large Language Models and thus achieve high data efficiency.
More Training Details . The detailed hyper-parameters of pre-training are listed in Tab. 11. Besides that, for image generation, we ignore the training loss of images which are the first element in the sequence. The text condition of the rest images are dropped with a 10% probability to improve classifier-free guidance sampling [32].
@@
Learning rate schedule cosine with warmup
@@
B.2 Supervised Fine-tuning
@@
Referring Expression Comprehension. Following previous works [10, 6], we train our model in the form of VQA, i.e., the prompt being Question: Provide the bounding box coordinate of the region this sentence describes: {object}, Answer: (x1,y1)(x2,y2). The generated bounding box is considered correct if its intersection over union (IoU) with the GT box is greater than 0.5. Only public available datasets, including datasets from RefCOCO [42], RefCOCO+ [58], and RefCOCOg [58] are utilized to train MM-Interleaved. See Tab. 13 for fine-tuning hyper-parameters.
@@
Visual Storytelling. Following previous works [43, 104, 66], MM-Interleaved is finetuned on the following three visual storytelling datasets respectively:
@@ B.3 Evaluation
LN-COCO [69] Text-to-image generation val-30k FID($\downarrow$[31]
@@
Task
Prompt Template
Zero-shot
Image Caption (short)
{image} a photo of
Image Caption (long)
{image} Please describe the image in detail. The image depicts
VQA (except VizWiz)
Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz QA
Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
Visual Dialog
@@
Supervised Fine-tuning
Image Caption (short)
{image} Provide a one-sentence caption for the provided image.
Image Caption (long)
{image} Please describe the image in detail.
VQA (except VizWiz)
Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz VQA
Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
Visual Dialog
@@
Appendix C Qualitative Results
C.1 VL Comprehension
@@ C.2 Image Generation
Appendix D Future Work
Future works of MM-Interleaved aims to further scale up both the model and data size, enabling end-to-end training with full-parameters from scratch. In addition, we believe that it is also worth exploring to build a comprehensive evaluation benchmark specifically designed for interleaved image-text modeling.

### v2

@@ 4.3 Ablation Study
(a) pre-training with 224 resolution
@@
(b) fine-tuning with 448 resolution
@@
Table 6: Ablation on using MMFS. “Generation” is text-to-image generation task. “ADE20k” is segmentation-to-image translation task. Others are text generation tasks. “# Token” indicates the number of input visual tokens for LLMs (32 by default).
@@
Table 7: Ablation on the design choice of MMFS. Different attention modules can be used in MMFS. We also ablate whether to add additional transition layer before feeding image features into MMFS. Dense cross-attention with Resampler transition on single scale and single image is similar to the cross attention used in Flamingo [3].
@@
Token Efficiency. As shown in Tab. 6(a), when equipped with MMFS, using only 32 visual tokens per image can outperform using 256 visual tokens without MMFS. Such results demonstrate the effectiveness of our method when the context length is limited. As shown in Tab. 6(b), the performance improvement of using MMFS becomes larger when increasing the input image resolution from 224 to 448. Such results indicate our method could better exploit the additional information gained from high resolution even with only 32 visual tokens.
@@
MMFS for Image Generation. Tab. 6(c) demonstrates the criticality of adding MMFS for segmentation-to-image translation task. This task is hard as it requires precise pixel-level information to align the given segmentation condition and image output properly. Without using MMFS, this task fails, showing extremely low mIoU results. Visualizations results in Appendix shows that the generated image without MMFS cannot preserve all spatial information, and the spatial alignment for the generated results is poor.
Comparison between Different Cross Attention Mechanisms. The ablation of adopting different cross-attention mechanisms for LLM is shown in Tab. 7. The overall performance drops when directly replacing MMFS with the vanilla dense cross-attention, possibly due to its slower convergence speed. We highlight that the model with Deformable attention performs significantly better than other attention mechanisms on TextVQA, indicating that Deformable attention could effectively and efficiently capture fine-grained information like text needed for the task, such as visual question answering in this case.
MMFS with Multi-Image and Multi-Scale. As shown in Tab. 8, adding multi-image and multi-scale for MMFS improves the performance. To better take advantage of the multi-image mechanism, we further evaluate our model with few-shot prompts following [3, 84]. As is illustrated in Fig. 6 (left), when the number of context images increases, MMFS continuously outperforms only using 32 visual tokens, and also benefits from further attending to the multi-scale feature maps of multiple images.
@@
Table 8: Ablation on multi-scale and multi-image usage of MMFS.
Figure 6: Left: Few-shot results on OKVQA and TextVQA. Right: Additional GFLOPs over using 32 visual tokens with different numbers of image and text inputs. $N_{v}$ and $k$ denote the number of visual-token per image and the number of few-shot examples, respectively. $N_{i}$ and $N_{t}$ denote the number of image and the number of subsequent text tokens for each image in the sequence. Note that $N_{v}=256$ does not support $N_{i}\geq$ 8 images due to 2048 LLM context length.
Computational Efficiency. Compared to only using 32 visual tokens, further integrating MMFS into LLM only increases by about 2%, 2%, 6% and 3% in FLOPs, #parameter, runtime and memory, respectively. When compared to using more visual tokens (i.e., 256) w/o MMFS, using MMFS with 32 visual tokens achieve better performance (see Tab. 6(a)) and is much more efficient, with 2.8$\times$ fewer FLOPs and 1.3$\times$ fewer runtime. Fig. 6 (right) shows the additional FLOPs over only using 32 visual tokens with synthesized interleaved image-text inputs. Our method consistently achieves lower or similar computation than using 256 visual tokens or using dense cross-attention.
5 Conclusion
This paper introduces MM-Interleaved, an end-to-end trained generative model for interleaved image-text comprehension and generation. It is built upon the Multi-Modal Feature Synchronizer (MMFS), which is proposed to enhance multi-modal LLMs by reducing the required visual tokens, enabling efficient extraction of visual details. Without using in-house data, our method achieves SOTA performance on various multi-modal benchmarks. We hope this work can contribute to the research on multi-modal LLMs for interleaved image-text scenarios.
Limitations. Both the quality and quantity of public interleaved image-text data are relatively low, making it difficult to exploit the full potential of interleaved generative modeling.
Negative impacts. As other multi-modal models, it may suffer from hallucination issues and potentially generate biased contents due to the noisy training data. We hope to invest more efforts to improve the quality and quantity of interleaved image-text data, with the aim of further improvements while ensuring its safety and reliability.
Appendix 0.A Implementation Details
0.A.1 Pre-training
@@
Data Concatenation Strategy. For image-text-pair datasets (i.e., LAION-2B, LAION-COCO [78]), we randomly sample multiple image-text pairs from the same dataset and concatenate them to the maximum context length (i.e., 2048) during pre-training. For interleaved image and text datasets (i.e., MMC4 [111]), we also split and concatenate the documents to form the training samples. Such concatenation strategy can utilize the full context window of Large Language Models and thus achieve high data efficiency.
More Training Details. The detailed hyper-parameters of pre-training are listed in Tab. 9. Besides that, for image generation, we ignore the training loss of images that are the first element in the sequence. The text condition of the rest images are dropped with a 10% probability to improve classifier-free guidance sampling [32].
@@
Learning rate schedule constant with warmup
@@
0.A.2 Supervised Fine-tuning
@@
Referring Expression Comprehension. Following previous works [10, 6], we train our model in the form of question answering, i.e., the prompt being Question: Provide the bounding box coordinate of the region this sentence
describes: {object}, Answer: (x1,y1)(x2,y2). The generated bounding box is considered correct if its intersection over union (IoU) with the GT box is greater than 0.5. Only public available datasets, including datasets from RefCOCO [42], RefCOCO+ [59], and RefCOCOg [59] are utilized to train MM-Interleaved. See Tab. 11 for fine-tuning hyper-parameters.
@@
Visual Storytelling. Following previous works [43, 106, 67], MM-Interleaved is finetuned on the following three visual storytelling datasets respectively. The finetuning hyper-parameters are listed in Tab. 13.
@@ 0.A.3 Evaluation
LN-COCO [70] Text-to-image generation val FID($\downarrow$[31]
@@
Task Prompt Template
Zero-shot Image Caption (short) {image} a photo of
Image Caption (long) {image} Please describe the image in detail. The image depicts
VQA (except VizWiz) Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz QA Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
@@
Supervised Fine-tuning Image Caption (short) {image} Provide a one-sentence caption for the provided image.
Image Caption (long) {image} Please describe the image in detail.
VQA (except VizWiz) Based on the image, please answer the question. {image}{question} Please provide an accurate answer within one word. The answer is:
VizWiz VQA Based on the image, please answer the question. {image}{question} When the provided information is insufficient, respond with ’Unanswerable’. Please provide an accurate answer within one word. The answer is:
@@
Appendix 0.B Additional Ablation Studies
Pre-training with Different Loss Terms. As shown in Tab. 16, MM-Interleaved achieves better performance when jointly trained with both $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ on comprehension and generation tasks, demonstrating the mutual benefits between the two loss terms. Moreover, setting $\lambda=10$ achieves a better balance between $\mathcal{L}_{NTP}$ and $\mathcal{L}_{NIP}$ empirically.
[…]
$\mathcal{L}_{NTP}+100~{}\mathcal{L}_{NIP}$ 106.2 31.1 29.8 24.5
$\mathcal{L}_{NTP}+10~{}\mathcal{L}_{NIP}$ 110.6 30.0 29.8 27.7
$\mathcal{L}_{NTP}+\mathcal{L}_{NIP}$ 110.0 31.4 29.3 26.0
$\mathcal{L}_{NTP}$ only 105.7 29.9 27.6
$\mathcal{L}_{NIP}$ only 34.2
Table 16: Pre-training with different loss terms.
[…]
110.6 30.0 29.8 27.7
107.0 32.2 28.7 22.5
102.7 32.0 27.3 22.0
Appendix 0.C Qualitative Results
0.C.1 Zero-shot Results
The zero-shot comprehension and generation visualizations of our method are shown in Figs. 7, 8, 9 and 10. Specifically, Fig. 7 demonstrates the emergent multi-modal in-context learning capability of our method, while Fig. 8 further illustrates the effectiveness of our method in understanding complex scenarios such as robotics, gaming and GUI interface. Fig. 9 exhibits that our method can also generate appropriate images based on the provided contexts of desired styles or concepts. Fig. 10 displays our method’s capability of generating images and texts interleavedly.
Figure 7: Zero-shot text generation with interleaved images and texts serving as multi-modal few-shot contexts.
Figure 8: Zero-shot text generation with interleaved images and texts on complex scenarios such as robotics, gaming and GUI interface.
Figure 9: Zero-shot image generation with interleaved images and texts. Our method can generate appropriate images based on the provided contexts of desired styles or concepts.
Figure 10: Zero-shot interleaved image and text generation for visual storytelling and multi-modal instructions.
0.C.2 Supervised Finetuning Results